At the beginning of the year, long-term debt of Century Company is $1,472,000 and total debt is $1,966,000. At the end of the year, long-term debt is $1,597,000 and total debt is $2,830,000. The interest paid is $92,000. What is the amount of the cash flow to creditors? -$33,000 $92,000 -$58,000 $67,000 $101,000
Answer #1

Net new long term debt=Ending long term debt-Beginning long term debt

=(1,597,000-1,472,000)=$125,000

Hence cash flow to creditors=Interest expense-Net new long term debt

=(92,000-125,000)

=($33,000)(Negative).